OpenStudy (anonymous):

what is equivalent to i^39?

OpenStudy (anonymous):

\(i^{3+36}=i^{3+4(9)}=i^{3}\cdot (i^{4})^9\)

OpenStudy (anonymous):

\(i^4=1\)

OpenStudy (anonymous):

you sir are wrong.

OpenStudy (anonymous):

lol, and can you back that up?

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

i'm waiting...

OpenStudy (anonymous):

go to the site and try that answer

OpenStudy (anonymous):

you, dear sir, are obviously unaware of how to simplify what was presented to you.

OpenStudy (anonymous):

i'm not here to hand out answers but you apparently are only here to collect such things. such a shame to waste an oppotunity to learn.

OpenStudy (anonymous):

ive just been on that problem for to long and im tired of trying to figure it out.

OpenStudy (anonymous):

to this point is correct: i^3 (i^4)^9 use these to finish it out: i^2 = -1 i^3 = -i i^4 = 1